OpenStudy (oh_hitz2523):

Solve the equation. x – = 3 A. x = –3 B. x = –2 C. x = 2 D. x = 3

OpenStudy (wckeller):

-x=3 is really -1*x=3 So what can you do to get X all alone on the left?

OpenStudy (oh_hitz2523):

the question is X - 2 over 7= 3 what is X

OpenStudy (wckeller):

OH

OpenStudy (wckeller):

Okay, so the equation is \[\frac{ x-2 }{ 7 } = 3\] First, how do we get rid of the fraction on the left?

OpenStudy (oh_hitz2523):

add?

OpenStudy (wckeller):

Unfortunately no... The operation to undo division is multiplication. We need to multiply the left and right side by 7.

OpenStudy (wckeller):

We would end up with: \[x-2=3*7\] \[3*7=21\] Therefore... \[x-2=21\]

OpenStudy (wckeller):

What can we do to get X all by itself?

OpenStudy (oh_hitz2523):

A. -3\[\frac{ 2 }{ 7 }\] B-2\[\frac{ 5 }{ 7 }\] C2\[\frac{ 5 }{ 7 }\] D3\[\frac{ 2 }{ 7 }\]

OpenStudy (oh_hitz2523):

@wckeller sorry i took so long

OpenStudy (oh_hitz2523):

those are the awsers

OpenStudy (wckeller):

Uhh.... Interesting.........

OpenStudy (wckeller):

The answer should be x=23 but that's nowhere near the answers on the list....

OpenStudy (oh_hitz2523):

the awsers are the numbers next to abcd plus the fractions

OpenStudy (wckeller):

So like A would be -3+(2/7) ?

OpenStudy (oh_hitz2523):

thank you so much

OpenStudy (will.h):

The correct answer is indeed x = 23 your options are wrong Report that question to your instructor you'd appear as a hero

OpenStudy (wckeller):

Thanks for the verification @Will.H I thought I was missing something.

OpenStudy (will.h):

Your welcome you may always check if your right by substituting the value you had in the position of the variable and see if the sides are equal

OpenStudy (oh_hitz2523):